Chem. If an ideal gas has a pressure of 2.13 atm, a temperature of 309 K, and has a volume of 59.73 L, how many moles of gas are in the sample A 6.8 increased its volume to 34.4 L. What was the initial volume of the balloon? s mol sample of freon gas was placed in a balloon. Adding 3.50 mol of freon gas to the balloon A hot air balloon is filled with 1.51x 106 L ofa 119 ℃, wha n ideal gas on a cool morning (11-C), The air is heated to ot is the volume of the air in the balloon after it is heated? Assume that none of the gas escapes from the balloon. The volume of a sample of hydrogen gas was decreased from 12.30 L to 5.59 Lat constant temperature. If the final pressure exerted by the hydrogen gas sample was 9.45 atm, what p gas exert before its volume was decreased? ressure did t he hydrogen

Please i need help to solve these problem
Answer #1

Solution:- (1) It's based on ideal gas law equation, PV = nRT

where, P is pressure in atm, V is volume in liters, n is moles of gas, R is universal gas constant and it's value is 0.0821 atm.L.mol-1.K-1

T is kelvin temperature.

We are asked to calculate the moles of the gas that is n.

n = PV/RT

let's plug in the given values..

n = (2.13 x 59.73)/(0.0821 x 309)

n = 5.01 mol

(2) According to Avogadro's law, At constant temperature and pressure, volume is directly proportional to the moles of the gas.

V1/V2 = n1/n2

V1 = V2*n1/n2

V1 = initial volume, V2 = final volume

n1 = initial moles and n2 = final moles

Given, n1 = 6.86 mol

n2 = 6.86 + 3.50 = 10.36 mol

V2 = 34.4 L

V1 = ?

On plugging in the values..

V1 = 6.86 x 34.4/10.36

V1 = 22.8 L

(3) It's based on Charles law. At constant pressure, volume of the gas is directly proportional to the kelvin temperature.

V1/T1 = V2/T2

Given, V1 = 1.51 x 106 L

T1 = 11 + 273 = 284 K

T2 = 119 + 273 = 392 K

V2 = ?

V2 = V1*T2/T1

V2 = 1.51 x 106 x 392/284

V2 = 2.08 x 106 L

(4) V1 = 12.30 L

V2 = 5.59 L

P2 = 9.45 atm

P1 = ?

It's Boyle's law. At constant temperature, volume of gas is inversely proportional to the pressure.

P1V1 = P2V2

P1 = P2V2/V1

P1 = 9.45 x 5.59/12.30

P1 = 4.29 atm